OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4). a.11 b. 19 c. 16 d. 8

OpenStudy (solomonzelman):

to find g(4), all you have to do is to plug in 4 for x. \[\huge\color{blue}{ g(4)=\huge\color{red}{ 4 } ^2+3 =16+3=19}\]SO, \[\huge\color{blue}{ g(4)=19}\]

OpenStudy (anonymous):

thank you i get it

OpenStudy (solomonzelman):

You welcome!

OpenStudy (anonymous):

can you help me with another question

OpenStudy (solomonzelman):

I can try.

OpenStudy (marigirl):

@SolomonZelman how did u type in an equation in colour?

OpenStudy (anonymous):

For the function f(x) = x + 7, what is the ordered pair for the point on the graph when x = 2b? A. (2b, 2b + 7) B. (2b, x + 7) C. (x, x + 7) D. (x, 2b + 7)

OpenStudy (solomonzelman):

\huge\color{ blue }{ } size color for example; \huge\color{ red }{ x } in equation editor that would look like \[\huge\color{ red }{ x } \]

OpenStudy (marigirl):

thanx :)

OpenStudy (anonymous):

Think of f(x) as being y. Thus if x =2b we know that is the value for x. Now what is the value for y. F(x) = x+7 or f(x) = 2b+7. Think of f(x) as being y. Answer would be A.